OpenStudy (anonymous):

Simplify: (3x^2 y^3)^2

OpenStudy (anonymous):

When you have an exponent on terms that are MULTIPLIED together, you can multiply each of the exponents of the terms inside the parentheses! So, \[ \left( 3x^{2}y^{3} \right)^2 = 3^2x^{2*2}y^{3*2}\] Note that this doesn't were for terms that are ADDED together. For example, \[ \left(3x^2+y^3 \right)^2 \neq 3^2x^{2*2}+y^{3*2}\] When you see addition, you have to FOIL Does this answer your question?

OpenStudy (anonymous):

Testing

OpenStudy (anonymous):

Sorry the replies weren't working on the safari browser

OpenStudy (anonymous):

Ok so for the exponent if (x^2)^2 is it 2+2 or 2x2 I guess in that seance either way is 4 but if it's x^3)^2 then wich would it be?

OpenStudy (anonymous):

You multiply them in this case. Because: \[ \left(x^3 \right)^2 = x^3x^3 = x^6 \]

OpenStudy (anonymous):

So if you see x's next to each other like \[ x^2x^3 \] then you'll want to add the exponents \[ x^2 x^3 = x^{2+3} = x^5 \] When you see exponents raise to a power, you multiply! \[ \left(x^2\right)^3 = x^{2*3} = x^6 \]

OpenStudy (anonymous):

Oh ok sorry that took me minute to catch I'm usually really good at math I'm just rusty since I haven't done math in 3 years and now I'm jumping back into grade 12u functions so now I'm trying to do the review questions and I'm stuck on all of them lol Ok so (3x^2y^3)^2 = 3x^4 y^6 is there any thing else I can do with this?

OpenStudy (anonymous):

haha no worries. \[ 3x^4y^6 \] is actually as far as you can simplify in this case

OpenStudy (anonymous):

Oops! you forgot to apply the exponent to the 3 too! it should be \[ 9x^4y^6 \]

OpenStudy (anonymous):

Ok thanks now to start a bunch of other questions lol